You have made a rectangular quilt that measures 4 feet by 5 feet. You have 10 square feet of fabric left, and decide to use it to add a border of uniform width around the quilt. What is the width of the border? Explain your answer.
Answer by fractalier(6550)

A couple of ways to explain this, so here's one way...
The original quilt is 4 x 5 feet...or 20 square feet...
Adding 10 more square feet brings the total to 30 square feet...
30 square feet can be looked at as 5 x 6 feet...
So what was 4 x 5 is now 5 x 6, meaning that 6 inches were added all around...
Not an algebraic way, but still easy to understand...
We could also solve
(4 + 2x)(5 + 2x) = (4)(5) + 10
which gives
x = 1/2 foot or 6 inches...